Make New Entry,
Make Followup Entry
User name Markowitz
Log entry time 08:15:16 on July 04, 2010
Entry number 329939
keyword=Owl shift summary: July 4, 2010
Owl Shift Summary: July 4, 2010
- Pete Markowitz
- James Beacham
- Victor Maxwell
- Lawrence Selvy
- 00:00 Hall is in controlled access for detector work. Bogdan,
Samvel, Sergey, Eric, and Albert are downstairs.
- 03:05 Crew in hall is done for now and ready to leave hall. Move
to empty target.
- 03:20 Turned on septum to 775.6 Amps, per MCC request.
- 03:30 Moved to Pb-D1 target. Request 0.5 uA beam. Start run 1636,
PS1=1, PS8=10, all others are 10000. T1 isLHRS S2m. Q1 of RHRS trips
off during run but allows us to reset from software. Since we are taking
left HRS data, it does not matter.
- 03:46 Start run 1638, PS2=1, LHRS S0.
- 03:53 Start run 1639, PS3=1, RHRS S2m. Beam tripped off so run is
- 04:00 Start run 1640, PS3=1, RHRS S2m.
- 04:05 Start run 1641, PS4=1, T4=T1&T3.
- 04:23 Start run 1642, PS6=1, T$&GC. Stopped to adjust PS.
- 04:30 Start run 1643, PS1-5,7=100000, PS6=1,PS8=100. Sergey sees a
coincidence peak in replay.
- 04:48 Start run 1644, Pedestal run.
- 04:51 Start run 1645, PS1-5,7=100000, PS6=1,PS8=100.
- 05:20 Access to fix missing channels and RHRS VDC. Phoned RC and
turned off septum.
- 06:00 Bogdan called from downstairs. He has removed the amplifier
from the RHRS for S0 since the signal was so strong and asked us to
change the HV to -1400 for both S0 PMTs on the RHRS.
- 06:17 Access finished, turning septum back on. Moved to empty
target to check beam location on dump viewer.
- 06:25 MCC verifies location on dump viewer is OK, and we move back
to Pb-D1 target. Ask for 0.5uA CW beam.
- 06:33 Start run 1650, PS1-5,7=100000, PS6=1,PS8=100. HV adjusted
- 07:13 Start run 1651, new HV. Noticed spot++ looks like we may be
hitting the edge of the frame. Stop run and decrease raster size from
12x8 to 10x6.
- 07:23 Start run 1652, raster size 10x6.
- 07:30 Spot still seems to show only half the parabola, so we are
asking MCC to move spot on target. Looking closely at the spot entry
from last night (329712)
it looks like we only have half of the range of their parabola, meaning
the spot is OK.
- 07:50 Bogdan asks us to call for an access.